OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 2x - y = 8. intercept = (2, 0); y-intercept = (0, -8) x-intercept = (4, 0); y-intercept = (0, -8) intercept = (2, 0); y-intercept = (0, 8) intercept = (4, 0); y-intercept = (0, 8)

OpenStudy (anonymous):

So y=mx+b 2x-y=8 -y=-2x+8 y=2x-8 X intercept: 2 Y intercept: -8